# Actively Learning Combinatorial Optimization Using a Membership Oracle

### v2

@@ Appendix B Instance Generation
For CSPP, we built two sets of instances inspired by the experimental setting of [1]. The first set of instances (MIT-Math) is designed according to the structure of the undergraduate courses in Mathematics at the Massachusetts Institute of Technology (MIT), whose list can be found here: http://student.mit.edu/catalog/m18a.html. Differently from [1], we took into account the entire list and we also considered every dependency inside other areas (Physics, Computer Science, Biology, etc.). For every course, the credits obtainable for passing the corresponding exam are specified as a triple $(a_{i},b_{i},c_{i})$, where the three values represent credits assigned for lectures and recitations, for laboratory, design, or field work, and for outside (independent) preparation, respectively. Each credit corresponds to 14 hours of activity, but while this equivalence is exact for $a_{i}$ and $b_{i}$, the hours actually dedicated to outside preparation vary from student to student. We operated few adjustments to the original data, mainly consisting in merging homologous courses, and we obtained a dataset with 150 courses whose structure of dependencies respect the definition of the college study plan problem. We generated 30 test instances, where the parameters related to each course $i$ are computed according to the following specifications from [1] ($B$ denotes the total hours budget equal to 1700, while $\mathcal{N}_{\text{T}}$ denotes the truncated normal distribution supported on $[0,B]$):
@@
    $\xi_{i}\sim\mathcal{N}_{\text{T}}(\mu_{i},\sigma_{i}^{2})/B$, where $\mu_{i}=14\cdot c_{i}$ and $\sigma_{i}\sim U[1,14×2]$;
@@
Appendix C Experimental Setups
Appendix D Additional Computational Results
We summarize in Table 1 different statistics obtained from the numerical experiments. In column “# Solved”, we report the number of instances where the exact optimal solution is found by the algorithm. In column “Ending Error (%)”, we report the relative objective error obtained when reaching the query limit. In column “Comput. Time (s)”, we report the total computational time required to solve the optimization subproblems in Algorithm 1. In column “# Queries”, we report the total number of queries. We use $\geq$" to indicate that more computational time and queries are needed to find the exact optimal solution of the instances. We observe that SEP+CUT solves the most instances in general, except for MIT-Math CSPP instances. The optimization subproblems required by the sampling strategy CUT seem more computationally demanding than the sampling strategy SIM.

### v3

@@ Appendix 0.C Instance Generation
For CSPP, we built two sets of instances inspired by the experimental setting of [1]. The first set of instances (MIT-Math) is designed according to the structure of the undergraduate courses in Mathematics at the Massachusetts Institute of Technology (MIT), whose list can be found here:
http://student.mit.edu/catalog/m18a.html.
Differently from [1], we took into account the entire list and we also considered every dependency inside other areas (Physics, Computer Science, Biology, etc.). For every course, the credits obtainable for passing the corresponding exam are specified as a triple $(a_{i},b_{i},c_{i})$, where the three values represent credits assigned for lectures and recitations, for laboratory, design, or field work, and for outside (independent) preparation, respectively. Each credit corresponds to 14 hours of activity, but while this equivalence is exact for $a_{i}$ and $b_{i}$, the hours actually dedicated to outside preparation vary from student to student. We operated few adjustments to the original data, mainly consisting in merging homologous courses, and we obtained a dataset with 150 courses whose structure of dependencies respect the definition of the college study plan problem. We generated 30 test instances, where the parameters related to each course $i$ are computed according to the following specifications from [1] ($B$ denotes the total hours budget equal to 1700, while $\mathcal{N}_{\text{T}}$ denotes the truncated normal distribution supported on $[0,B]$):
@@
    $\xi_{i}\sim\mathcal{N}_{\text{T}}(\mu_{i},\sigma_{i}^{2})/B$, where $\mu_{i}=14\cdot c_{i}$ and $\sigma_{i}\sim U[1,14\texttimes 2]$;
@@
Appendix 0.D Additional Computational Results
We summarize in Table 1 different statistics obtained from the numerical experiments.
@@ Appendix 0.D Additional Computational Results
In column “# Solved”, we report the number of instances where the exact optimal solution is found by the algorithm. In column “Ending Error (%)”, we report the relative objective error obtained when reaching the query limit. In column “Comput. Time (s)”, we report the total computational time required to solve the optimization sub-problems in Algorithm 1. In column “# Queries”, we report the total number of queries. We use $\geq$" to indicate that more computational time and queries are needed to find the exact optimal solution of the instances. We observe that SEP+CUT solves the most instances in general, except for MIT-Math CSPP instances. The optimization sub-problems required by the sampling strategy CUT seem more computationally demanding than the sampling strategy SIM.